OpenStudy (anonymous):

x^3 + x^2 - 2x - 30 = 0 what are the solutions?

OpenStudy (anonymous):

x = 3, x = -2-i sqrt(6), x = -2+i sqrt(6)

OpenStudy (anonymous):

Well, at first I guessed for the 3, then divided the polynomial by (x-3), then I used the quadratic formula.